\begin{thebibliography}{00}

\bibitem{Bambusi_A} D. Bambusi and G. Gaeta,
 On persistence of invariant tori and a theorem by Nekhoroshev,
 {\it MPEJ} {\bf 8} (2002), 1.

\bibitem{Dazord_A} P. Dazord and T. Delzant,
 Le probl\`{e}me g\'{e}n\'{e}ral des variables actions-angles,
 {\it J. Diff. Geom.} {\bf 26} (1987), 223--251.

\bibitem{Duistermaat_A} J. J. Duistermaat,
 On global action-angle coordinates,
 {\it Comm. Pure Appl. Math.} {\bf 33} (1980), 687--706.

\bibitem{Fasso_A} F. Fass\`{o},
 Superintegrable Hamiltonian systems: geometry and perturbations,
 {\it Acta Appl. Math.} {\bf 87} (2005), 93--121.

\bibitem{Fasso2_A} F. Fass\`{o} and T. Ratiu,
 Compatibility of symplectic structures adapted to noncommutatively
 integrable systems,
 {\it J. Geom. Phys.} {\bf 27} (1998), 220.

\bibitem{Flaschka_A} H. Flaschka,
 A remark on integrable Hamiltonian systems,
 {\it Phys. Lett. A} {\bf 131} (1988), 505--508.

\bibitem{Fiorani1_A} E. Fiorani, G. Giachetta and G. Sardanashvily,
 Geometric quantization of time-dependent completely integrable
 Hamiltonian systems,
 {\it J. Math. Phys.} {\bf 43} (2002), 5013--5025.

\bibitem{Fiorani2_A} E. Fiorani, G. Giachetta and G. Sardanashvily,
 An extension of the Liouville-Arnold theorem for the noncompact
 case,
 {\it Il Nuovo Cimento} {\bf B118} (2003), 307--317.

\bibitem{Fiorani3_A} E. Fiorani and G. Sardanashvily,
 Noncommutative integrability on noncompact invariant manifolds,
 {\it J. Phys. A: Math. Gen.} {\bf 39} (2006), 14035--14042.

\bibitem{Fiorani4_A} E. Fiorani and G. Sardanashvily,
 Global action-angle coordinates for completely integrable systems
 with noncompact invariant submanifolds,
 {\it J. Math. Phys.} {\bf 48} (2007), 32901.

\bibitem{Gaeta_A} G. Gaeta,
 The Poincar\'{e}-Lyapunov-Nekhoroshev theorem,
 {\it Ann. Phys.} {\bf 297} (2002), 157--173.

\bibitem{Giachetta_A} G. Giachetta, L. Mangiarotti and G. Sardanashvily,
 Bi-Hamiltonian partially integrable systems,
 {\it J. Math. Phys.} {\bf 44} (2003), 1984--1997.

\bibitem{Meigniez_A} G. Meigniez,
 Submersions, fibrations and bundles,
 {\it Trans. Am. Math. Soc.} {\bf 354} (2002), 3771--3787.

\bibitem{Mishchenko_A} A. S. Mishchenko and A. T. Fomenko,
 Generalized Liouville method of integration of Hamiltonian systems,
 {\it Funct. Anal. Appl.} {\bf 12} (1978), 113--121.

\bibitem{Nekhoroshev2_A} N. N. Nekhoroshev,
 Action-angle variables and their generalization,
 {\it Trans. Moscow Math. Soc.} {\bf 26} (1972), 180--198.

\bibitem{Nekhoroshev_A} N. N. Nekhoroshev,
 The Poincar\'{e}-Lyapunov-Liouvulle-Arnold theorem,
 {\it Funct. Anal. Appl.} {\bf 28} (1994), 128--129.

\bibitem{Palais_A} R. S. Palais,
 A global formulation of the Lie theory of transformation groups,
 {\it Mem. Am. Math. Soc.} {\bf 22} (1994).

\bibitem{Arnold_B} V. I. Arnold (ed.),
 {\it Dynamical Systems III}
 (Springer, Berlin, 1988).

\bibitem{Giachetta_B} G. Giachetta, L. Mangiarotti and G. Sardanashvily,
 {\it New Lagrangian and Hamiltonian Methods in Field Theory}
 (World Scientific, Singapore, 1997).

\bibitem{Greub_B} W. Greub, S. Halperin and R. Vanstone,
 {\it Connections, Curvature and Cohomology (Vol.2)}
 (Academic Press, New York, 1973).

\bibitem{Griffiths_B} P. Griffiths and J. Harris,
 {\it Principles of Algebraic Geometry}
 (John Wiley \& Sons, New York, 1978).

\bibitem{Guillemin_B} V. Guillemin and S. Sternberg,
 {\it Symplectic Techniques in Physics}
 (Cambridge University Press, Cambridge, 1984).

\bibitem{Kobayashi_B} S. Kobayashi and K. Nomizu,
 {\it Foundations of Differential Geometry (Vol. 1)}
 (Interscience Publishers, New York, 1969).

\bibitem{Libermann_B} P. Libermann and C. M. Marle,
 {\it Symplectic Geometry and Analytical Mechanics}
 (Reidel, Dordrecht, 1987).

\bibitem{Woodhouse_B} N.M.J. Woodhouse,
 {\it Geometric quantization (2nd ed.)}
 (Clarendon Press, Oxford, 1994).

\end{thebibliography}

\documentclass{article}
\usepackage{amsthm}

\theoremstyle{definition}
\newtheorem{definition}{Definition}
\theoremstyle{plain}
\newtheorem{theorem}{Theorem}
\theoremstyle{remark}
\newtheorem{remark}{Remark}
\theoremstyle{plain}
\newtheorem{lemma}{Lemma}

\begin{document}

\begin{center}

\begin{Large}
\textbf{Geometrical aspects of integrable systems}
\end{Large}
\bigskip

\begin{large}
Emanuele Fiorani
\end{large}

Department of Mathematics and Informatics, University of Camerino \\
62032 Camerino (MC), Italy \\
emanuele.fiorani@unicam.it

\end{center}
\bigskip

\begin{abstract}
We review some basic theorems on integrability of Hamiltonian
systems, namely the Liouville-Arnold theorem on complete
integrability, the Nekhoroshev theorem on partial integrability and
the Mi\-sh\-che\-nko-Fo\-me\-nko theorem on noncommutative
integrability, and for each of them we give a version suitable for
the noncompact case. We give a possible global version of the
previous local results, under certain topological hypotheses on the
base space. It turns out that locally affine structures arise
naturally in this setting.
\bigskip

\noindent \textbf{Keywords:} Integrable Hamiltonian systems; global
action-angle coordinates; locally affine structures.
\end{abstract}

\section{Introduction}

There are some well known theorems treating integrable Hamiltonian
systems on a $2n$-dimensional symplectic manifold $(Z, \Omega )$:
the so called Liouville-Arnold theorem (or
Li\-ou\-vi\-lle-Mi\-ne\-ur-Arno\-ld theorem) for completely
integrable systems (see e.g. \cite{Arnold_B,Libermann_B}), the
Nekhoroshev theorem (or Poin\-car\'{e}-Lya\-pou\-nov-Nek\-horo\-shev
theorem) for partially integrable systems (see
\cite{Nekhoroshev_A}), and the Mi\-sh\-che\-nko-Fo\-me\-nko theorem
for noncommutative (i.e. non-Abelian) systems (see
\cite{Fasso_A,Mishchenko_A}). All of them state the existence of a
neighborhood of a compact invariant submanifold (which is a torus
$\mathbf{T}^n$) which is topologically trivial (i.e. a product), and
the existence of particular adapted coordinates on it, called
action-angle coordinates. These theorems have been extended to the
case of noncompact invariant submanifold (which is a cylinder
$\mathbf{R}^{n-h} \times \mathbf{T}^h$); see
\cite{Fiorani1_A,Fiorani2_A,Fiorani3_A,Giachetta_A}. Sections 2, 3
and 4 intend to compare the compact and noncompact situations.

If invariant submanifolds are compact, topological obstructions to
the global triviality of the whole $Z$ and to the global existence
of action-angle coordinates have been studied (see
\cite{Dazord_A,Duistermaat_A,Nekhoroshev2_A}). Here we aim to extend
this study to the case of noncompact invariant submanifolds; a
possible approach is shown in Section 5 and follows that of
\cite{Fiorani4_A}. The main idea is to provide a sufficient
condition in order that a principal bundle with Abelian structure
group be trivial.

Section 6 is devoted to the interplay between integrability and
locally affine structures. The motivation is to clarify the meaning
of a certain connection naturally arising on the invariant
submanifolds; see for instance \cite{Fasso2_A,Woodhouse_B}. For this
purpose it is needed to study in some details the relation between
flat and parallelizable manifolds and, in this context, the role of
the torsion of a flat connection is considered. Some peculiarity of
spheres and Euclidean spaces are outlined.

Throughout this paper all manifolds are smooth and paracompact and
all maps and structures are smooth, $(Z, \Omega)$ is a
$2n$-dimensional connected symplectic manifold and $(C^{\infty}(Z),
\{ \})$ is the Poisson algebra of (real) functions on $Z$. Given a
function $F: Z \rightarrow \mathbf{R}$, its Hamiltonian vector field
$\Omega^{\sharp}$d$F$ is indicated with $X_F$.

\section{Complete integrability}

Perhaps Theorem \ref{ComplComp} below is the prototype of
integrability theorems; it treats Hamiltonian systems with the
maximum possible number of (independent) functions in involution.
This theorem, together with further discussions, can be found for
instance in the books \cite{Arnold_B,Libermann_B} which are standard
references.

\begin{definition}
A subset $(F_1, \ldots, F_n)$ of $(C^{\infty}(Z), \{ \})$ is called
a completely integrable system on $Z$ if at every point of $Z$
$\mathrm{d}F_1, \ldots, \mathrm{d}F_n$ are linearly independent and
$\{F_i, F_j\} = 0$ $\forall \, i,j = 1, \ldots , n$.
\end{definition}

\begin{theorem} \label{ComplComp}
 Let us assume the following:
 \begin{itemize}
  \item[(i)] $F = (F_1, \ldots, F_n): Z \rightarrow \mathbf{R}^n$ is a
  completely integrable system;

  \item[(ii)] There is a compact component $M$ of a fiber of $F$.
 \end{itemize}

 Then:

 \begin{itemize}
  \item[(I)] There exists a (saturated) neighborhood $U$ of $M$ such that
  $F|_U$ is a trivial principal bundle with structure group the torus
  $\mathbf{T}^n$, and $M$ is one of its fibers;

  \item[(II)] Given standard coordinates $(\varphi^{\lambda})$ on
  $\mathbf{T}^n$, the neighborhood $U$ is provided with Darboux
  coordinates $(I_{\lambda}, \varphi^{\lambda})$ such that
  $$
  \Omega = \mathrm{d}I_{\lambda} \wedge \mathrm{d}\varphi^{\lambda}
  $$
  and $F_{\lambda} = F_{\lambda}(I_{\mu})$ on $U$.
 \end{itemize}
\end{theorem}

\begin{remark}
\begin{itemize}
 \item[(i)] If $F: A \rightarrow B$ is a submersion, the compact components
 of the fibers of $F$ make up an {\em open} subset of $A$; see the
 discussion on page 358 of \cite{Guillemin_B}.

 \item[(ii)] A submersion $F: A \rightarrow B$ ($B$ connected) with
 compact fibers having constant (finite) number of components is a
 bundle; indeed it can be shown that, under these hypotheses, $F$ is
 a proper map. Alternatively, if the fibers of $F$ are diffeomorphic
 to a Euclidean space $\mathbf{R}^n$ then $F$ is a bundle.
 See \cite{Meigniez_A}.

 Thus (i) and (ii) can be used to give a simple proof
 of Theorem \ref{ComplComp}.

 \item[(iii)] In contrast with (ii), a submersion with fibers
 diffeomorphic to a product $\mathbf{R}^n \times K$, where $K$ is a
 compact manifold, need {\em not} be a bundle. See \cite{Giachetta_B},
 Example 1.2.2, in which a submersion onto $\mathbf{R}$ with fibers
 all diffeomorphic to $\mathbf{R} \times \mathbf{S}^1$ is shown
 to have no trivial neighborhood of 0 $\in \mathbf{R}$.
\end{itemize}
\end{remark}

It is possible to remove the compactness hypothesis for $M$ in the
previous theorem in the following way; see
\cite{Fiorani1_A,Fiorani2_A} for a proof and further discussions.

\begin{theorem} \label{ComplNonComp}
 Let us assume the following:
  \begin{itemize}
  \item[(i)] $F = (F_1, \ldots, F_n): Z \rightarrow \mathbf{R}^n$ is a
  completely integrable system;

  \item[(ii)] The fibers of $F$ are connected and mutually diffeomorphic;

  \item[(iii)] Each Hamiltonian vector field $X_{F_{\lambda}}$ is complete.
 \end{itemize}

Then:

 \begin{itemize}
  \item[(I)] $F$ is a bundle with typical fiber the cylinder
  \begin{equation}
   \mathbf{R}^{n-h} \times \mathbf{T}^{h}; \label{CylCompl}
  \end{equation}
  moreover there is a covering of $F(Z)$ with simply connected domains
  $\{V_{\alpha}\}$ such that  over each $V_{\alpha}$ $F$ is
  a principal bundle with structure group (\ref{CylCompl});

  \item[(II)] Given standard coordinates $(y^{\lambda})$ on (\ref{CylCompl}),
  each fiber of $F$ has a (saturated) neighborhood $U$ provided with
  Darboux coordinates $(I_{\lambda}, y^{\lambda})$ such that
  $$
  \Omega = \mathrm{d}I_{\lambda} \wedge \mathrm{d}y^{\lambda}
  $$
  and $F_{\lambda} = F_{\lambda}(I_{\mu})$ on $U$.
 \end{itemize}
\end{theorem}

\begin{remark}
\begin{itemize}
 \item[(i)] If we remove the completeness hypothesis (iii) in Theorem
 \ref{ComplNonComp}, then the fibers of $F$ can be much more complicated than
 $\mathbf{R}^{n-h} \times \mathbf{T}^{h}$: see \cite{Flaschka_A}
 where it is shown that if we take for $F$ a polynomial with two complex
 variables, say $(z_1, z_2) \in \mathbf{C}^2 \simeq \mathbf{R}^4$, then the
 functions Re$(F)$ and Im$(F)$ are in involution with respect to a
 certain symplectic structure on $\mathbf{R}^4$. Thus we can have
 generic Riemann surfaces as fibers of such an $F$.

 \item[(ii)] If one of the fibers of $F$ is compact, then it must be a torus
 and by (i) of Remark 3 there is a saturated neighborhood of it made up
 of such compact fibers; moreover each $X_{F_{\lambda}}$ is complete on them.
 Thus all the hypotheses of Theorem \ref{ComplNonComp} are verified and we
 can see Theorem \ref{ComplComp} as a special case of Theorem
 \ref{ComplNonComp}.
\end{itemize}
\end{remark}

\section{Partial integrability}

Theorem \ref{ParComp} below treats the case of a Hamiltonian system
with a number of (independent) functions in involution lesser than
the maximum possible. It first appeared in the article of N. N.
Nekhoroshev \cite{Nekhoroshev_A} but a more detailed proof of it,
together with discussions and some worked out examples, can be found
in \cite{Bambusi_A,Gaeta_A}. Moreover, Proposition 27.8 of
\cite{Guillemin_B} is a related result: it gives a sufficient
condition for a foliation with a compact leaf to be a trivial bundle
around that leaf.

\begin{definition}
Let $k$ be an integer, $1 \leq k \leq n$.
A subset $(F_1, \ldots, F_k)$ of $(C^{\infty}(Z), \{ \})$ is called
a partially integrable system on $Z$ if at every point of $Z$
$\mathrm{d}F_1, \ldots, \mathrm{d}F_k$ are linearly independent and
$\{F_i, F_j\} = 0$ $\forall \, i,j = 1, \ldots , k$.
\end{definition}

\begin{theorem} \label{ParComp}
 Let us assume the following:
 \begin{itemize}
  \item[(i)] $F = (F_1, \ldots, F_k): Z \rightarrow \mathbf{R}^k$ is a
  partially integrable system;

  \item[(ii)] There is a compact leaf $M$ of the foliation spanned by the
  Hamiltonian vector fields $X_{F_1}, \ldots, X_{F_k}$;

  \item[(iii)] There is a closed curve $\gamma$ on $M$ such that
  D$\Psi_{\gamma}$ does not have 1 as an eigenvalue, where
  $\Psi_{\gamma}$ is the Poincar\'{e} map corresponding to $\gamma$.
 \end{itemize}

Then:

 \begin{itemize}
  \item[(I)] There exists a {\em submanifold} $N$ of $Z$, $N$ of dimension 2$k$
  and $N \supseteq M$, such that $F|_N$ is a trivial principal bundle with
  structure group $\mathbf{T}^n$, and $M$ is one of its fibers;

  \item[(II)] Given standard coordinates $(\varphi^{\lambda})$ on $\mathbf{T}^k$,
  $N$ is provided with Darboux coordinates $(I_{\lambda},
  \varphi^{\lambda})$ such that
  $$
  \Omega = \mathrm{d}I_{\lambda} \wedge \mathrm{d}\varphi^{\lambda}
  $$
  and $F_{\lambda} = F_{\lambda}(I_{\mu})$ on $N$.
 \end{itemize}
\end{theorem}

It is possible to remove the compactness hypothesis for the leaf $M$
in the previous theorem in the following way; see
\cite{Fiorani2_A,Giachetta_A} for a proof and further discussions.

\begin{theorem} \label{ParNonComp}
 Let us assume the following:
 \begin{itemize}
  \item[(i)] $F = (F_1, \ldots, F_k): Z \rightarrow \mathbf{R}^k$ is a
  partially integrable system;

  \item[(ii)] The leaves of the foliation spanned by the Hamiltonian
  vector fields $X_{F_1}, \ldots, X_{F_k}$ are mutually
  diffeomorphic and are the fibers of a submersion
  $\mathcal{F}: Z \rightarrow B$;

  \item[(iii)] Each Hamiltonian vector field $X_{F_{\lambda}}$ is complete.

\end{itemize}

Then:

 \begin{itemize}
  \item[(I)] $\mathcal{F}$ is a bundle with typical fiber the cylinder
  \begin{equation}
   \mathbf{R}^{k-h} \times \mathbf{T}^{h}; \label{CylPar}
  \end{equation}
  moreover there is a covering of $\mathcal{F}(Z)$ with simply connected
  domains $\{V_{\alpha}\}$ such that  over each $V_{\alpha}$ $F$ is
  a principal bundle with structure group (\ref{CylPar});

  \item[(II)] Given standard coordinates $(y^{\lambda})$ on (\ref{CylPar}),
  each fiber of $\mathcal{F}$ has a (saturated) neighborhood $U$
  provided with Darboux coordinates
  $(I_{\lambda}, p_A, q^A, y^{\lambda})$ such that
  $$
   \Omega = \mathrm{d}I_{\lambda} \wedge \mathrm{d}y^{\lambda}
    + \mathrm{d}p_A \wedge \mathrm{d}q^A
  $$
  and $F_{\lambda} = F_{\lambda}(I_{\mu})$ on $U$.
 \end{itemize}
\end{theorem}

\begin{remark}
\begin{itemize}
 \item[(i)] If $k = n$ this theorem reduces to Theorem
 \ref{ComplNonComp} about complete integrability.

 \item[(ii)] The Hamiltonian vector fields $\{X_{I_{\lambda}}
 = \frac{\partial}{\partial y^{\lambda}}\}$ are obviously complete,
 but the Hamiltonian vector fields $\{X_{p_A} = \frac{\partial}{\partial q^A}\}$
 and $\{X_{q^A} = -\frac{\partial}{\partial p_A}\}$ may fail to be
 complete.

 If $\{X_{p_A}\}$ are complete, then the functions $(I_{\lambda}, p_A)$
 verifies the hypotheses of Theorem \ref{ComplNonComp} about
 complete integrability on $U$. Moreover if also $\{X_{q^A}\}$ are complete,
 then the functions $(I_{\lambda}, p_A, q^A)$ verifies the hypotheses
 of Theorem \ref{SupNonComp} below about noncommutative integrability on $U$.

 This is indeed more than it is needed; see part (ii) of Remark
 \ref{RemSup}.
\end{itemize}
\end{remark}

\section{Noncommutative integrability}

In the previous sections we considered subsets of $(C^{\infty}(Z),
\{\})$ of at most $n$ (independent) functions which commute, i.e.
generate a commutative Lie algebra. Now we want to treat the case of
a number of functions greater that $n$; of course in this case they
cannot be mutually commuting (if they are independent).

The following theorem goes in this direction; it first appeared in
the the article of A. S. Mishchenko and A. T. Fomenko
\cite{Mishchenko_A} in the special case in which the functions $F_1,
\ldots, F_k$ generate a certain \emph{finite} dimensional Lie
algebra, in general not commutative (here $n \leq h < 2n$). Theorem
\ref{SupComp} in the present form (which does not need this
restriction) can be found in \cite{Fasso_A}.

\begin{theorem} \label{SupComp}
 Let us assume the following:
 \begin{itemize}
  \item[(i)] $F = (F_1, \ldots, F_k): Z \rightarrow \mathbf{R}^k$ is a
  submersion with compact and connected fibers (here $n \leq k < 2n$);

  \item[(ii)] $\{F_{\lambda}, F_ {\mu}\} = s_{\lambda\mu}(F)$ with the
  $k \! \times \! k$ matrix $(s_{\lambda\mu})$ of constant rank $\, 2(k-n)$.
 \end{itemize}

 Then:

 \begin{itemize}
  \item[(I)] $F$ is a bundle with typical fiber the torus
  $\mathbf{T}^{2n-k}$; moreover there is a covering of $F(Z)$ with
  simply connected domains $\{V_{\alpha}\}$ such that over each
  $V_{\alpha}$ $F$ is a principal bundle with structure group
  $\mathbf{T}^{2n-k}$;

  \item[(II)] Given standard coordinates $(\varphi^i)$ on
  $\mathbf{T}^{2n-k}$, each fiber of $F$ has a (saturated)
  neighborhood $U$ provided with Darboux coordinates
  $(I_i, p_A, q^A, \varphi^i)$ such that
  $$
  \Omega = \mathrm{d}I_i \wedge \mathrm{d}\varphi^i
   + \mathrm{d}p_A \wedge \mathrm{d}q^A
  $$
  and $F_{\lambda} = F_{\lambda}(I_i, p_A, q^A)$ on $U$.

  \noindent Given a Hamiltonian function $\mathcal{H}$ of the
  system, it depends only on coordinates $(I_i)$ on $U$.
 \end{itemize}
\end{theorem}

It is possible to remove the compactness hypothesis for the fibers
of $F$ in the previous theorem in the following way; the proof given
here is an improvement of the proof that can be found in
\cite{Fiorani3_A} together with further discussions and a worked out
example.

\begin{theorem} \label{SupNonComp}
 Let us assume the following:
 \begin{itemize}
  \item[(i)] $F = (F_1, \ldots, F_k): Z \rightarrow \mathbf{R}^k$ is a
  submersion with connected and mutually diffeomorphic fibers
  (here $n \leq k < 2n$);

  \item[(ii)] $\{F_{\lambda}, F_ {\mu}\} = s_{\lambda\mu}(F)$ with the
  $k \! \times \! k$ matrix $(s_{\lambda\mu})$ of constant rank $\, 2(k-n)$.

  \item[(iii)] Each Hamiltonian vector field $X_{F_{\lambda}}$ is complete.
 \end{itemize}

 Then:

 \begin{itemize}
  \item[(I)] $F$ is a bundle with typical fiber the cylinder
  \begin{equation}
   \mathbf{R}^{2n-k-h} \times \mathbf{T}^{h}; \label{CylNonComm}
  \end{equation}
  moreover there is a covering of $F(Z)$ with simply connected
  domains $\{V_{\alpha}\}$ such that over each $V_{\alpha}$ $F$ is a
  principal bundle with structure group (\ref{CylNonComm}).

  \item[(II)] Given standard coordinates $(y^i)$ on
  (\ref{CylNonComm}), each fiber of $F$ has
  a (saturated) neighborhood $U$ provided with Darboux coordinates
  $(I_i, p_A, q^A, y^i)$ such that
  $$
  \Omega = \mathrm{d}I_i \wedge \mathrm{d} y^i
   + \mathrm{d}p_A \wedge \mathrm{d}q^A
  $$
  and $F_{\lambda} = F_{\lambda}(I_i, p_A, q^A)$ on $U$.

  \noindent Given a Hamiltonian function $\mathcal{H}$ of the
  system, it depends only on coordinates $(I_i)$ on $U$.
 \end{itemize}
\end{theorem}

\begin{proof}
$F$ is a submersion, so $F(Z)$ is an open subset of $\mathbf{R}^k$
and we can choose standard coordinates $(x_{\lambda})$ on it.

Since $\forall \: \lambda, \mu = 1, \ldots k$ we have
$\{F_{\lambda}, F_{\mu}\} = s_{\lambda\mu}(F)$, it is possible to
provide $F(Z)$ with a Poisson structure $\{\}_F$ whose Poisson
tensor is just
\begin{equation} \label{Poisson}
s_{\lambda\mu}(x) \frac{\partial}{\partial x_{\lambda}} \wedge
\frac{\partial}{\partial x_{\mu}}
\end{equation}
and $F$ is a Poisson morphism.

Since $\forall \, x \! \in \! F(Z)$ we have rank$(s_{\lambda\mu}(x))
= 2(k-n)$, around each $x \in F(Z)$ there are $2n-k$ independent
Casimir functions $\{C_i\}$ of $\{\}_F$ such that the Hamiltonian
vector fields $\{X_{F^*C_i}\}$ span the fibers of $F$. See
\cite{Libermann_B} for these facts, where $\{\}_F$ is called
\emph{coinduced Poisson structure}.

Now if we consider the Hamiltonian vector field
$\{X_{F_{\lambda}}\}$ instead of the functions $\{F_{\lambda}\}$,
from (ii) it follows that
 $$
[X_{F_{\lambda}}, X_{F_{\mu}}] = - \frac{\partial
s_{\lambda\mu}}{\partial x_{\alpha}} X_{F_{\alpha}}
 $$
i.e. on each fiber of $F$ they generate a \emph{finite} dimensional
Lie algebra; then each linear combination of $\{X_{F_{\lambda}}\}$
is complete; see \cite{Palais_A}. In particular, each Hamiltonian
vector field of the functions $\{F^*C_i\}$
\begin{equation} \label{LinCombFields}
X_{F^*C_i} = \frac{\partial C_i}{\partial x_{\alpha}} X_{F_{\alpha}}
\end{equation}
is complete.

Then it is possible to apply the previous Theorem \ref{ParNonComp}
to the partial integrable system on $Z$ given by $\{F^*C_i\}$.
\end{proof}

\begin{remark} \label{RemSup}
\begin{itemize}
 \item[(i)] If $k = n$ this theorem reduces to Theorem
 \ref{ComplNonComp} about complete integrability.

 \item[(ii)] The Hamiltonian vector fields $\{X_{I_i}
 = \frac{\partial}{\partial y^i}\}$ are obviously complete,
 thus the functions $(I_i)$ verifies the hypotheses of Theorem
 \ref{ParNonComp} about partial integrability on $U$ if we take
 as $\mathcal{F}$ the projection $(I_i, p_A, q^A, y^i)
 \mapsto (I_i, p_A, q^A)$.

 The Hamiltonian vector fields $\{X_{p_A} = \frac{\partial}{\partial q^A}\}$
 and $\{X_{q^A} = -\frac{\partial}{\partial p_A}\}$ may fail to be
 complete. If $\{X_{p_A}\}$ are complete, then the functions
 $(I_i, p_A)$ verifies the hypotheses of Theorem \ref{ComplNonComp}
 about complete integrability on $U$.

 In any case, the completeness hypothesis (iii) here is used only to
 guarantee that each field $X_{F^*C_i}$ in (\ref{LinCombFields}) is
 complete.
\end{itemize}
\end{remark}

\section{Integrability and global action-angle coordinates}

So far all the theorems stated the existence of action-angle
coordinates on a (saturated) neighborhood of any invariant manifold.
It is natural now to investigate the global picture: for the compact
case standard references are the articles of P. Dazord and T.
Delzant \cite{Dazord_A}, J. J. Duistermaat \cite{Duistermaat_A} and
N. N. Nekhoroshev \cite{Nekhoroshev2_A}. In what follows there is a
possible way to treat the case of noncompact invariant manifolds;
Theorem \ref{GlobPar}, \ref{GlobSup} and \ref{GlobCompl} can be
found in \cite{Fiorani4_A}; let us note that Lemma \ref{LemmaGlob}
and Theorem \ref{PrincGlob} here establish a basic sufficient
condition for a principal bundle with Abelian structure group to be
trivial, and permit to simplify the proofs of the theorems contained
in \cite{Fiorani4_A}.

\begin{lemma} \label{LemmaGlob}
Let $P \rightarrow B$ be a principal bundle and let us assume that
its structure group is a product $G_1 \times G_2$.

Then it is canonically isomorphic to the product principal bundle
$P_1 \times P_2 \rightarrow B$ where $P_1 \rightarrow B$ is the
quotient principal bundle $P/G_2 \rightarrow B$ and $P_2 \rightarrow
B$ is $P/G_1 \rightarrow B$.
\end{lemma}

\begin{proof}
Using the fact that we have canonical projections of $G_1 \times
G_2$ onto its factors $G_1$ and $G_2$, the proof is a
straightforward check.
\end{proof}

\begin{theorem} \label{PrincGlob}
Let $P \rightarrow B$ be a principal bundle and let us assume the
following:
\begin{itemize}
 \item[(i)] The structure group $G$ is Abelian;

 \item[(ii)] The base $B$ verifies $H^2(B, \textbf{Z})=0$.
\end{itemize}
Then the principal bundle $P \rightarrow B$ is trivial.
\end{theorem}

\begin{proof}
Since the structure group $G$ is Abelian, it is isomorphic to
$\textbf{R}^n$ or to the cylinder $\textbf{R}^{n-h} \times \textbf{T}^h$,
where $n =\,$dim$G$ and $h \geq 1$. In the first case, the result
follows immediately from the existence of a global section;
see \cite{Kobayashi_B}.

In the second case, it is possible to reduce the group
$\textbf{R}^{n-h} \times \textbf{T}^h$ to its maximal compact
subgroup $\textbf{T}^h$; then, by Lemma \ref{LemmaGlob}, it is
isomorphic to the product of $h$ copies of principal $U(1)$-bundles
over the same base $B$. Since $H^2(B, \textbf{Z})=0$, each of these
bundles has trivial first Chern class and consequently it is
trivial; see \cite{Griffiths_B}. Thus the original principal bundle
is trivial.
\end{proof}

The following theorems are global versions of Theorem
\ref{ParNonComp} and \ref{SupNonComp} about partial and
noncommutative integrability.

\begin{theorem} \label{GlobPar}
 Let us assume the following:
 \begin{itemize}
  \item[(i)] $F = (F_1, \ldots, F_k): Z \rightarrow \mathbf{R}^k$ is a
  partially integrable system;

  \item[(ii)] The leaves of the foliation spanned by the Hamiltonian
  vector fields $X_{F_1}, \ldots, X_{F_k}$ are mutually
  diffeomorphic and are the fibers of a submersion
  $\mathcal{F}: Z \rightarrow B$;

  \item[(iii)] Each Hamiltonian vector field $X_{F_{\lambda}}$ is
  complete;

  \item[(iv)] The base $B$ is simply connected and $H^2(B,
  \mathbf{Z}) = 0$.
\end{itemize}

Then:

 \begin{itemize}
  \item[(I)] $\mathcal{F}$ is a trivial principal bundle with structure group
  the cylinder $\mathbf{R}^{k-h} \times \mathbf{T}^{h}$;

  \item[(II)] Given standard coordinates $(y^{\lambda})$ on
  $\mathbf{R}^{k-h} \times \mathbf{T}^{h}$, $Z$ is provided with
  coordinates $(I_{\lambda}, x^A, y^{\lambda})$ such that
  $$
   \Omega = \mathrm{d}I_{\lambda} \wedge \mathrm{d}y^{\lambda}
    + \Omega^{\lambda}_A \mathrm{d}I_{\lambda} \wedge \mathrm{d}x^A
    + \Omega_{AB} \mathrm{d}x^A \wedge \mathrm{d}x^B
  $$
  and $F_{\lambda} = F_{\lambda}(I_{\mu})$ on $Z$.
 \end{itemize}
\end{theorem}

\begin{theorem} \label{GlobSup}
 Let us assume the following:
 \begin{itemize}
  \item[(i)] $F = (F_1, \ldots, F_k): Z \rightarrow \mathbf{R}^k$ is a
  submersion with connected and mutually diffeomorphic fibers
  (here $n \leq k < 2n$);

  \item[(ii)] $\{F_{\lambda}, F_ {\mu}\} = s_{\lambda\mu}(F)$ with the
  $k \! \times \! k$ matrix $(s_{\lambda\mu})$ of constant rank $\, 2(k-n)$.

  \item[(iii)] Each Hamiltonian vector field $X_{F_{\lambda}}$ is
  complete;

  \item[(iv)] There is an open subset $V$ of $F(Z)$ admitting $2n-k$
  independent Casimir functions of the coinduced Poisson structure on $F(Z)$
  as in the proof of Theorem \ref{SupNonComp} which is simply connected
  and such that $H^2(V, \mathbf{Z}) = 0$.
 \end{itemize}

Then:

 \begin{itemize}
  \item[(I)] The restriction $F|V: F^{-1}(V) \rightarrow V$ is a trivial
  principal bundle with structure group the cylinder
  $\mathbf{R}^{2n-k-h} \times \mathbf{T}^{h}$;

  \item[(II)] Given standard coordinates $(y^i)$ on
  $\mathbf{R}^{2n-k-h} \times \mathbf{T}^{h}$, $F^{-1}(V)$ is provided with
  coordinates $(I_{\lambda}, x^A, y^{\lambda})$ such that
  $$
   \Omega = \mathrm{d}I_{\lambda} \wedge \mathrm{d}y^{\lambda}
    + \Omega_{AB} \mathrm{d}x^A \wedge \mathrm{d}x^B
  $$
  and $F_{\lambda} = F_{\lambda}(I_{\mu}, x^A)$ on $F^{-1}(V)$.

  \noindent Given a Hamiltonian function $\mathcal{H}$ of the
  system, it depends only on coordinates $(I_{\lambda})$ on $F^{-1}(V)$.
 \end{itemize}
\end{theorem}

\begin{remark} The proof of Theorem \ref{SupNonComp} shows that
each $x \in F(Z)$ has a neighborhood $V$ admitting $2n-k$
independent Casimir functions; thus the result holds over any such
$V$ which is simply connected and such that $H^2(V, \mathbf{Z}) =
0$.
\end{remark}

The following theorem is a global version of Theorem
\ref{ComplNonComp}; note that in this case the set of hypotheses is
simpler than in the previous cases in that it requires only
topological properties on the image $F(Z)$ of the original
submersion $F$.

\begin{theorem} \label{GlobCompl}
 Let us assume the following:
 \begin{itemize}
  \item[(i)] $F = (F_1, \ldots, F_n): Z \rightarrow \mathbf{R}^n$ is a
  completely integrable system;

  \item[(ii)] The fibers of $F$ are connected and mutually diffeomorphic;

  \item[(iii)] Each Hamiltonian vector field $X_{F_{\lambda}}$ is
  complete;

  \item[(iv)] $F(Z)$ is simply connected and $H^2(F(Z),
  \mathbf{Z}) = 0$.
 \end{itemize}

 Then:

 \begin{itemize}

  \item[(I)] $F$ is a {\rm trivial} principal bundle with structure group
  the cylinder $\mathbf{R}^{n-h} \times \mathbf{T}^{h}$;

  \item[(II)] Given standard coordinates $(y^{\lambda})$ on
  $\mathbf{R}^{n-h} \times \mathbf{T}^{h}$, $Z$ is provided
  with Darboux coordinates $(I_{\lambda}, y^{\lambda})$ such that
  $$
  \Omega = \mathrm{d}I_{\lambda} \wedge \mathrm{d}y^{\lambda}
  $$
  and $F_{\lambda} = F_{\lambda}(I_{\mu})$ on $Z$.
 \end{itemize}
\end{theorem}

\begin{remark} Since in hypothesis (iv) we assume $F(Z)$
to be simply connected, $H^2(F(Z), \mathbf{Z}) = 0$ is equivalent to
$\pi_2(F(Z)) = 0$, i.e. (iv) can be restated by requiring $F(Z)$ to
be 2-connected. The same holds for the base manifold $B$ in Theorem
\ref{GlobPar} and for the neighborhood $V$ in Theorem \ref{GlobSup}.
\end{remark}

\section{Integrability and locally affine structures}

In what follows $M$ is a connected and $m$-dimensional manifold; we
recall the following two standard definitions.

\begin{definition}
Let $\nabla$ be a linear connection on $M$ and let $R_{\nabla}$,
$T_{\nabla}$ be its curvature and torsion tensors respectively;
\begin{itemize}
 \item $\nabla$ is flat when $R_{\nabla} = 0$;

 \item $\nabla$ is a locally affine structure on $M$ when $R_{\nabla} =
 0$ and $T_{\nabla} = 0$.
\end{itemize}
\end{definition}

\begin{definition}
Let $\{X_1, \dots, X_m\}$ be a parallelization of $M$ and let $X$,
$Y$ be vector fields on $M$, $Y = b^jX_j$ with $b^j$ global smooth
functions on $M$. The expression
\begin{equation}
\nabla_XY = X(b^j)X_j \label{ConnPar}
\end{equation}
is easily seen to define a linear connection on $M$, called the
connection of the parallelization.
\end{definition}

The basic properties of the connection (\ref{ConnPar}) are
summarized in the following theorem.

\begin{theorem} \label{ProprConnPar}
Let $\{X_1, \dots, X_m\}$ be a parallelization of $M$ and let
$\nabla$ be as in (\ref{ConnPar}). Then:
\begin{itemize}
 \item[(I)] $\nabla$ is a flat linear connection;

 \item[(II)] $T_{\nabla}(X_i, X_j) = -[X_i, X_j]$;

 \item[(III)] Each $X_i$ is a geodesic vector fields for $\nabla$,
 i.e. each integral curve of $X_i$ is a geodesic of $\nabla$;

 \item[(IV)] Given another parallelization $\{\bar{X}_1, \ldots, \bar{X}_m\}$ of
 $M$, it defines the same $\nabla$ if and only if  $\bar{X}_j = G^i_jX_i$
 $\forall \ j = 1, \ldots, m$, with $(G^i_j)$ some constant
 invertible matrix.
\end{itemize}
\end{theorem}

\begin{proof}
Statement (I) can be found in \cite{Greub_B}, but they all follow
from direct computations. In particular, to prove (IV) it is
sufficient to consider two parallelizations $\{X_i\}$ and
$\{\bar{X}_j\}$ of $M$ and observe that there must be a relation
$\bar{X}_j = G^i_j(x)X_i$ between the two, with $(G^i_j(x))$ an
invertible matrix of smooth functions on $M$. Then the condition
that their associated connections $\nabla$ and $\bar{\nabla}$ must
be equal leads to the equivalent condition $X_k(G^i_j(x)) = 0$
$\forall i,j,k = 1, \ldots, m$ and $\forall x \in M$; the facts that
$\{X_i\}$ is a basis at every point and that $M$ is connected yields
that the matrix $(G^i_j)$ must be constant.
\end{proof}

The following theorem states that when $M$ is simply connected also
the converse of Theorem \ref{ProprConnPar} holds.

\begin{theorem} \label{ConvProprConnPar}
Let us assume the following:
\begin{itemize}
 \item[(i)] There is a flat linear connection $\nabla$ on $M$;

 \item[(ii)] $M$ is simply connected.
\end{itemize}
Then there is a parallelization $\{X_1, \ldots, X_m\}$ of $M$ such
that its connection is $\nabla$.
\end{theorem}

\begin{proof}
The proof can be found in \cite{Greub_B}, but also a direct proof
can be given. Let us fix a point $x_0 \in M$ and a basis $\{X_1,
\ldots, X_m\}$ of $T_{x_0}M$. Since $\nabla$ is flat and $M$ simply
connected, it is possible, through parallel transport, to move this
basis to any point $x \in M$ getting a basis of $T_xM$ and the
result does not depend on the path followed. Thus we get a
parallelization of $M$ still indicated with $\{X_1, \ldots, X_m\}$;
it can be seen that it is smooth and $\nabla_{X_i} X_j = 0$ $\forall
\ i, j = 1, \ldots, m$. It follows that, given any vector field $Y =
b^jX_j$ on $M$ with $b^j$ global smooth functions on $M$, we get for
the original connection $\nabla$ the formula $\nabla_{X_i}Y =
X_i(b^j)X_j$, which coincides with the formula (\ref{ConnPar}) for
the connection of the parallelization $\{X_1, \ldots, X_m\}$.
\end{proof}

\begin{remark} Let us consider the following relation on the set of
all the parallelizations of a manifold $M$: two parallelizations
$\{X_1\, \ldots, X_m\}$ and $\{\bar{X}_1, \ldots, \bar{X}_m\}$ are
equivalent if and only if there is a constant invertible matrix
$(G^i_j)$ such that
\begin{equation} \label{Parallelization}
\bar{X}_j = G^i_jX_i \quad \forall j = 1, \ldots, m.
\end{equation}
It is clear that (\ref{Parallelization}) is an equivalence relation.
From Theorem \ref{ProprConnPar} and \ref{ConvProprConnPar} it
follows that on a \emph{simply connected} manifold $M$, flat linear
connections $\nabla$ are in 1-to-1 correspondence with equivalence
classes of (\ref{Parallelization}).
\end{remark}

We can improve Theorem \ref{ProprConnPar} and \ref{ConvProprConnPar}
by considering the torsion of $\nabla$.

\begin{theorem} \label{ImprProprConnPar}
Let $\{X_1, \ldots, X_m\}$ be a parallelization of $M$ with
commuting vector fields and let $\nabla$ be as in (\ref{ConnPar});
then $\nabla$ is a locally affine structure on $M$.

Conversely, let $\nabla$ be a locally affine structure on $M$ simply
connected; then there is a parallelization $\{X_1, \ldots, X_m\}$ of
$M$ with commuting vector fields such that its connection is
$\nabla$.
\end{theorem}

\begin{proof}
It all follows from Theorem \ref{ProprConnPar} and
\ref{ConvProprConnPar}, remembering the formula $T_{\nabla}(X_i,
X_j) = -[X_i, X_j]$.
\end{proof}

\begin{remark} \label{RemCompl}
\begin{itemize}

 \item[(i)] Let us consider $\{X_1, \ldots, X_m\}$ a parallelization of $M$
 with commuting and {\em complete} vector fields: in this case,  besides
 the locally affine structure (\ref{ConnPar}), we can obtain information
 about $M$. Indeed, the composition $\Phi$ of the flows of the vector fields
 $\{X_1, \ldots, X_m\}$ define a transitive action of $\mathbf{R}^m$
 on $M$. Thus $M$ must be diffeomorphic to $\mathbf{R}^m$ modulo the
 discrete isotropy group of one of its point, which yields
 $\mathbf{R}^{m-h} \times \mathbf{T}^h$ for some $h \in \{0, \ldots,
 m\}$. See for instance \cite{Arnold_B} for this standard argument.
 This situation occurred in in particular in all the theorems of the
 preceding sections.

 \item[(ii)] Of course, also nonorientable manifolds can admit a
 locally affine structure: examples are the M\"{o}bius strip
 and the Klein bottle; they obviously are not simply connected.

\end{itemize}
\end{remark}

It is known that among all the spheres $\mathbf{S}^m$, the
parallelizable ones are $\mathbf{S}^1$, $\mathbf{S}^3$ and
$\mathbf{S}^7$. Putting together the facts of this section, we
obtain the following additional properties; we consider only the
case $\mathbf{S}^m$ with $m \geq 2$, the case $\mathbf{S}^1$ being
trivial.

\begin{theorem} \label{Spheres}
Let us consider the spheres $\mathbf{S}^m$, $m \geq 2$.
\begin{itemize}
 \item[(I)] The only spheres admitting a flat linear connection are
 $\mathbf{S}^3$ and $\mathbf{S}^7$;

 \item[(II)] No sphere $\mathbf{S}^m$ admits a locally
 affine structure.
\end{itemize}
\end{theorem}

\begin{proof} $\mathbf{S}^3$ and $\mathbf{S}^7$ are parallelizable,
thus they admit a flat linear connection. Moreover, each
$\mathbf{S}^m$, $m \geq 2$, is simply connected but for $m \neq 3,
7$ \ they are not parallelizable; thus for $m \neq 3, 7$ \ they
cannot admit a flat linear connection. This proves (I).

According to (I), there remain only $\mathbf{S}^3$ and
$\mathbf{S}^7$ to be checked. Since they are simply connected, if
they admitted a locally affine structure they would be
parallelizable with commuting vector fields. Since they are compact,
these vector fields would be complete and by (i) of Remark
\ref{RemCompl} they would be diffeomorphic to $\mathbf{T}^3$ and
$\mathbf{T}^7$ respectively, which is impossible because they are
not simply connected. This proves (II).
\end{proof}

We can also obtain the following version of the Cartan-Hadamard
theorem:

\begin{theorem} \label{CartanAdamard}
Let us assume the following:
\begin{itemize}
 \item[(i)] There is a locally affine structure $\nabla$ on $M$;

 \item[(ii)] $\nabla$ is geodesically complete, i.e. all of
 its geodesics can be defined on $\mathbf{R}$;

 \item[(iii)] $M$ is simply connected.
\end{itemize}
Then there is a diffeomorphism $\Psi: M \rightarrow \mathbf{R}^m$
such that $\nabla$ corresponds under $\Psi$ to the standard locally
affine structure of \ $\mathbf{R}^m$.
\end{theorem}

\begin{proof}
By Theorem \ref{ImprProprConnPar} $M$ is parallelizable with
commuting vector fields, which are complete by (III) of Theorem
\ref{ProprConnPar}; then, after we fix a starting point $\alpha \in
M$, the map $\Phi: \mathbf{R}^m \rightarrow M$ of (i) of Remark
\ref{RemCompl} must define a diffeomorphism because $M$ is simply
connected. Moreover a computation shows that if $\Phi: t \mapsto p$
then
$$
T_t\Phi: \mbox{e}_i \longmapsto X_i(p) \quad \forall \ i=1, \dots, m
$$
where $(\mbox{e}_1, \dots, \mbox{e}_m)$ is the standard basis of
$\mathbf{R}^m$. Thus it is enough to take as $\Psi$ the function
$\Phi^{-1}$.
\end{proof}

\begin{remark}
\begin{itemize}
 \item[(i)] More generally, we can state the following: no compact
 and simply connected manifold admits a locally affine structure;
 the proof is exactly the same as that of part (II) of Theorem
 \ref{Spheres}. In particular, there we implicitly considered the
 standard differentiable structure of $\mathbf{S}^7$, but the result
 of part (II) holds for all its exotic differentiable structures.

 \item[(ii)] From Theorem \ref{CartanAdamard} it follows that
 on $\mathbf{R}^m$,  after a diffeomorphism, all the geodesically complete
 locally affine structures can be brought to the standard one.
 In particular, since there we implicitly considered the standard
 differentiable structure for $\mathbf{R}^4$, no exotic
 differentiable structure on $\mathbf{R}^4$ admits a
 geodesically complete locally affine structure.
\end{itemize}
\end{remark}

It is well known that locally affine structures arise naturally with
certain foliations of a symplectic manifold. We recall only the
basic definition and construction; for further discussions we refer
to \cite{Dazord_A,Fasso_A,Fasso2_A,Woodhouse_B} and references
therein for the following terminology and contruction.

\begin{definition}
Let $\mathcal{F}$ be a foliation of the symplectic manifold $(Z, \Omega)$.

The polar foliation of $\mathcal{F}$ is, if it exists, the foliation
$\mathcal{F}^{\perp}$ of $Z$ with the property that the tangent spaces
of its leaves are the symplectic complements of the tangent spaces
of the leaves of $\mathcal{F}$.

A foliation $\mathcal{F}$ which admits a polar foliation
$\mathcal{F}^{\perp}$ is called symplectically complete.
\end{definition}

If we have a symplectically complete foliation $\mathcal{F}$ of $Z$
with isotropic leaves, then these leaves admit a locally affine
structure, which is given by
\begin{equation} \label{OmegaComm}
\nabla^{\Omega}_XY = \Omega^{\sharp} L_X \Omega^{\flat}(Y)
\end{equation}
where $X$, $Y$ are vector fields tangent to the leaves of $\mathcal{F}$.

\begin{remark} \label{RemFibers}
It is readily observed that the fibers of the following submersions
\begin{itemize}
 \item[(i)] $F$ of Theorem \ref{ComplNonComp}

 \item[(ii)] $\mathcal{F}$ of Theorem \ref{ParNonComp}

 \item[(iii)] $F$ of Theorem \ref{SupNonComp}
\end{itemize}
define foliations of $Z$ which are isotropic and symplectically complete,
so they admit the locally affine structure $\nabla^{\Omega}$
defined in (\ref{OmegaComm})
\end{remark}

It is natural to investigate the nature of the connection
$\nabla^{\Omega}$ for the cases listed in Remark \ref{RemFibers};
the following theorem states that it is nothing else than the
connection of appropriate parallelizations.

\begin{theorem}
Let us consider the submersions (i), (ii) and (iii) of Remark
\ref{RemFibers} and the connection $\nabla^{\Omega}$ defined on
their fibers by (\ref{OmegaComm}); $\nabla^{\Omega}$ coincides with
the connection $\nabla$ (\ref{ConnPar}) of the following
parallelizations:
\begin{itemize}
 \item For (i) the parallelization given by the
 Hamiltonian vector fields $\{X_{F_1}, \ldots, X_{F_n}\}$;

 \item For (ii) the parallelization given by the
 Hamiltonian vector fields $\{X_{F_1}, \ldots, X_{F_k}\}$;

 \item For (iii) the parallelization given by the
 Hamiltonian vector fields $\{X_{F^*C_1}, \ldots, X_{F^*C_{2n-k}}\}$
 found at the end of the proof of Theorem \ref{SupNonComp}.
\end{itemize}
\end{theorem}

\begin{proof}
To simplify the notations, let us call generically $X_i$ the
Hamiltonian vector field of the function $F_i$; then this proof will
adapt to all the three previous statements. Given a vertical vector
$Y = b^jX_j$, we have
\begin{eqnarray*}
 \nabla^{\Omega}_{X_k}Y & = & \Omega^{\sharp}L_{X_k}\Omega^{\flat}(b^jX_j) =
 \Omega^{\sharp}L_{X_k}(b^j\mbox{d}F_j) = \\
 & = & \Omega^{\sharp}[i_{X_k}\mbox{d}(b^j\mbox{d}F_j)] =
 \Omega^{\sharp}[i_{X_k}(\mbox{d}b^j \wedge \mbox{d}F_j)] = \\
 & = & \Omega^{\sharp}[X_k(b^j)\mbox{d}F_j] = X_k(b^j)X_j = \\
 & = & \nabla_{X_k}Y
\end{eqnarray*}
where we have used that in our case $\Omega(X_i, X_j) = \{F_i, F_j\}
= 0$.
\end{proof}

\end{document}